OpenStudy (anonymous):

.76M KOH how do i find the pH?

OpenStudy (anonymous):

pOH = -log(OH-) in this case since KOH is a strong base, hence KOH will dissociates completely with water to form OH-. pOH = -log(0.76) pH+pOH=14 pH= 14- pOH.